ETHANE SIMMETRIC STRECHING VIBRATIONAL NORMAL MODES

\underline{MARIA VILLA}, Departamento de Quimica,,Universidad Autonoma Metropolitana, Av. San Rafael Atlixco, Col. Vicentina, Iztapalapa, Mexico D.F. 09340, MEXICO; M. LUISA SENENT, Departamento de Astrofisica Molecular e Infrarroja, Instituto de Estructura de la Materia, C.S.I.C., Serrano 113B, Madrid 28006, SPAIN; and RICARDO HIDALGO, Departamento de Quimica,,Universidad Autonoma Metropolitana, Av. San Rafael Atlixco, Col. Vicentina, Iztapalapa, Mexico D.F. 09340, MEXICO.

Ethane is reported as a non-rigid molecule taking as variable the θ methyl internal rotational angle. In this work the symmetric C-H stretching and the C-C stretching normal modes of vibration will be considered as function of the internal rotational angle. The potential function is expressed as a Taylor expansion in terms of the normal modes, and the expansion variables are evaluated at different θ angle and finally they are expressed and fitted to a θ function. It is quite important to emphasize the symmetry of the normal modes of vibration and the non-rigid group before writing a properly simetrized potential function.